A trailer is pulling a car with a force of 900N on a level road, if the car is moving at a velocity of 70km/h.
How much work is done by the trailer on the car in 15 minutes.
1
Expert's answer
2020-08-10T19:54:13-0400

"W=F\cdot l=F\cdot vt=900(N)\cdot19.44(m/s)\cdot900(s)=15.75\cdot10^6(J)"